# LLM Output Homogenization is Task Dependent

### v1

@@ 6.1 Evaluation Datasets
We sample 300 total prompts from the following datasets to use in evaluation of output homogenization. These datasets were chosen to achieve coverage across our task taxonomy (c.f. Table 3). For random sampling, we first shuffle the dataset using a random seed of 38, then select the required number of prompts in order from the shuffled dataset.
  • Community Alignment (Zhang et al. (2025a)): A diverse human preference dataset containing user-generated prompts. We use 50 randomly-sampled prompts from the subset of user-generated first-turn prompts in English. Users were instructed to “ask, request, or talk to the model about something important to you or that represents your values. This could be related to work, religion, family, relationships, politics, or culture.”
  • MATH-500 (Lightman et al. (2023)): A subset of the MATH dataset Hendrycks et al. (2021). We use 10 randomly-sampled prompts from each of the 5 difficulty levels.
  • NoveltyBench (Zhang et al. (2025b)): A dataset of creative tasks where multiple distinct and high-quality outputs are expected. We use their entire curated dataset of 100 prompts.
  • SimpleQA (Wei et al. (2024)): A dataset of short, fact-seeking queries across diverse topics. The prompts were created to be challenging for frontier models (e.g. GPT-4o accuracy $40\%$). We use 50 randomly-sampled prompts.
  • WildBench (Lin et al. (2025)): A subset of the WildChat dataset Zhao et al. (2024). WildChat is a corpus of 1 million user-ChatGPT conversations. WildBench is a filtered subset of WildChat such that tasks are diverse and challenging for models. We use 50 randomly-sampled prompts from the WildBench-V2.
@@ 6.2 Task Classification Into Our Taxonomy
We classify prompts into our taxonomy using the following judge prompt. Note that we exclude category E (Problem Solving or Design Subjective) because we did not identify any datasets containing prompts for this category. In all our analysis, we analyze prompts based on the task category chosen by majority vote across the three models (GPT-4o, Claude-4-Sonnet, and Gemini-2.5-Flash). Thus, our experiments focus on evaluating models’ ability to promote diversity when given the task category. We do not evaluate models’ ability to correctly categorize tasks. However, all three models agree on the categorization for 79% of prompts, while two of the three models agree on an additional 19%. We use the GPT-4o label for the remaining cases (<<2%).
Read the prompt below and decide which task category it belongs to. Only output a single category letter (A, B, C, D, E, F, or G) without any additional text. For prompts that have objective responses, choose from categories A, B, C, or D. For prompts that have subjective responses, choose from categories E, F, or G.
@@
E - Encyclopedia Inquiry Subjective: Task to generate information about real-world societies, traditions, events, or social domains, where there are credible references.
F - Creative Generation Subjective: Task to generate a response that involves creative expression where there are potentially infinite subjective responses.
G - Advice or Opinion Subjective: Task to generate a response that gives advice, opinions, or feedback on specific topics or scenarios.
Table 3: Number of prompts per dataset and taxonomy category.
Community Math 500 Novelty Bench Simple QA Wild Bench Total
Alignment
Well-Specified Objective 4 13 1 50 3 71
Underspecified Objective 0 0 32 0 3 35
Random Generation 0 0 11 0 0 11
Problem-Solving Objective 1 37 0 0 12 50
Encyclopedia Inquiry 11 0 3 0 5 19
Creative Writing 1 0 25 0 24 50
Advice or Opinion 33 0 28 0 3 64
Total 50 50 100 50 50 300
@@ 6.3 Sampling Strategies
We generate 5 responses per prompt using temperature sampling, in-context regeneration, and system prompt sampling. For in-context regeneration, the first response is generated with temperature 1.0 for GPT-4o and Gemini-2.5-Flash, and temperature 0.5 for Claude-4-Sonnet. Subsequent responses are generated using the regeneration prompts below. For system prompt sampling, multiple responses are extracted from a single generation. For all sampling strategies, we set the nucleus sampling parameter to 0.9 and the maximum number of output tokens to 1024.
@@
We use the following prompt for the general system prompting approach. This prompt is similar to the one used in Zhang et al. (2025a), except they use the phrase “Generate {num_responses} that represent diverse values”. We generalize this instruction to apply to all the categories in our taxonomy, since Zhang et al. (2025a) focus on more subjective tasks.
@@
Zhang et al. (2025b) use the following prompt for in-context regeneration. We call this a “general” prompt because there is no task dependence.
@@ 6.4 Measuring Functional Diversity
We validate the LLM-judges on a stratified random sample of 225 response pairs across models, tasks, and sampling strategies. Two authors independently labeled these prompts for functional diversity. Table 5 reports the inter-rater agreement between annotators and each LLM-judge. When compared to the ensemble of LLM-judges, annotator 1 achieved 77.7% agreement and annotator 2 achieved 79.0% agreement. This is comparable to Zhang et al. (2025b)’s evaluation of functional diversity in creative tasks, where a fine-tuned judge achieved 79% agreement with human annotators.
@@
Table 5: Inter-Rater Agreement Between Human Annotators & LLM-Judges
@@
Category Task-Anchored Functional Diversity Definition Diversity Judge Options

### v2

@@ 6 Additional Experiment Details
6.1 Taxonomy Crosswalks
Our taxonomy is grounded in existing literature on LLM output homogenization and diversity. Specifically, we observe that many studies evaluate homogenization in specific task domains, suggesting that problematic notions of homogenization are task dependent. We designed our taxonomy to cover a variety of these task categories. Table 3 maps our taxonomy to related works that have evaluated output homogenization for each task category. To our knowledge, discussions of response variance in well-specified objective prompts (category A) are often found in studies of confabulation, not in the homogenization literature.
We further show that our task taxonomy captures many real-world LLM use cases identified in recent work. chatterji2025people create a task taxonomy based on ChatGPT usage, which we map to our taxonomy in Table 4. Similarly, tamkin2024clio provide a list of common tasks based on Claude usage, which we map to our taxonomy in Table 5. We show that all real-world task categories in these previous works map to at least one category in our taxonomy (for text-based tasks). Many real-world task categories appear to correspond with multiple task categories in our taxonomy. For example, the “Practical Guidance” category in chatterji2025people may correspond to Problem-Solving Subjective (Category E) or Advice/Opinions (Category H). This illustrates how our task categories are meant to capture different functional diversity concepts, while the categories in these other works are meant to summarize usage trends. For “Practical Guidance” tasks, we would consider them as Problem-Solving Subjective if responses span different partially verifiable solutions, and as Advice/Opinions if responses span different non-verifiable perspectives or views.
Ultimately, our task categories represent one categorization of functional diversity concepts, and a task may fall outside our taxonomy or correspond to multiple categories. In these cases, the model may resume its default behavior instead of using our task-anchored sampling technique, or promote diversity based on one of the applicable categories. Our approach is further generalizable to alternative taxonomies or task categories.
Table 3: Crosswalk of Our Taxonomy and Previous Output Homogenization Studies
Task Category Previous Works (Non-Exhaustive)
A. Well-Specified Singular Objective wei2024simpleQA
B. Underspecified Singular Objective zhang2025noveltybench
C. Random Generation hopkins2023can; zhang2025noveltybench
D. Problem-Solving Objective lee2024enhancing; slocumdiverse; wu2025generative
E. Problem-Solving or Design Subjective ma2024exploring; yang2025prompts
F. Encyclopedia Inquiry sharma2024generative; sui2025critical; wright2025epistemic
G. Creative Writing anderson2024homogenization; doshi2024generative; lanchantin2025diverse; moon2024homogenizing; moon2025impersonal; padmakumar2024does; wu2025generative; zhang2025noveltybench
H. Advice or Opinions agarwal2025ai; durmus2023towards; santurkar2023whose; shahid2025llms; zhang2025cultivating
Table 4: Crosswalk of Task Categories for ChatGPT Usage with Our Taxonomy
ChatGPT Task Category Categories in Our Taxonomy
c.f. Table 3 in chatterji2025people
Writing (Edit or Critique Provided Text, Personal Writing or Communication, Translation, Argument or Summary Generation, Write Fiction) Underspecified Objective (B),
Creative Writing (G),
Advice or Opinions (H)
Practical Guidance (How-To Advice, Tutoring or Teaching, Creative Ideation, Health, Fitness, Beauty, or Self-Care) Problem-Solving Subjective (E),
Advice or Opinions (H)
Technical Help (Mathematical Calculation, Data Analysis, Computer Programming) Problem-Solving Objective (D),
Problem-Solving Subjective (E)
Multimedia (Create an Image, Analyze an Image, Generate or Retrieve Other Media) N/A (Our taxonomy is limited to text-based tasks)
Seeking Information (Specific Info, Purchasable Products, Cooking & Recipes) Well-Specified Objective (A),
Underspecified Objective (B),
Encyclopedia Inquiry (F),
Advice or Opinions (H)
Self-Expression (Greetings & Chitchat, Relationships & Personal Reflection, Games & Role Play) Creative Writing (G),
Advice or Opinions (H)
Table 5: Crosswalk of Top 10 Task Categories in Claude Usage with Our Taxonomy
Claude Task Category Categories in Our Taxonomy
c.f. Figure 6 in tamkin2024clio
Web and mobile application
development assistance
Problem-Solving Objective (D),
Problem-Solving Subjective (E)
Content creation and communication
assistance across disciplines
Creative Writing (G)
Multidisciplinary academic research
and writing assistance
Well-Specified Objective (A),
Underspecified Objective (B),
Encyclopedia Inquiry (E)
Education and career development
assistance
Advice or Opinions (H)
Implement and optimize diverse
AI/ML technologies and applications
Problem-Solving Objective (D),
Problem-Solving Subjective (E)
Business strategy and operations
assistance across industries
Problem-Solving Subjective (E),
Advice or Opinions (H)
Multilingual NLP, translation, and
linguistic analysis services
Underspecified Objective (B),
Creative Writing (G)
DevOps and cloud infrastructure
implementation and troubleshooting
Problem-Solving Objective (D),
Problem-Solving Subjective (E)
Digital marketing and SEO
optimization assistance
Problem-Solving Subjective (E),
Advice or Opinions (H)
Data analysis, visualization, and
management assistance
Problem-Solving Subjective (E),
Advice or Opinions (H)
@@ 6.2 Evaluation Datasets
We sample 350 total prompts from the following datasets to use in evaluation of output homogenization. These datasets were chosen to achieve coverage across our task taxonomy (c.f. Table 6). For random sampling, we first shuffle the dataset using a random seed of 38, then select the required number of prompts in order from the shuffled dataset.
  • Community Alignment (zhang2025cultivating): A diverse human preference dataset containing user-generated prompts. We use 50 randomly-sampled prompts from the subset of user-generated first-turn prompts in English. Users were instructed to “ask, request, or talk to the model about something important to you or that represents your values. This could be related to work, religion, family, relationships, politics, or culture.”
  • MacGyver (tian2024macgyver): A dataset of creative problem-solving tasks. We use 50 randomly-sampled prompts from the subset of “solvable” problems that require “unconventional” solutions.
  • MATH-500 (lightman2023let): A subset of the MATH dataset hendrycks2021measuring. We use 10 randomly-sampled prompts from each of the 5 difficulty levels.
  • NoveltyBench (zhang2025noveltybench): A dataset of creative tasks where multiple distinct and high-quality outputs are expected. We use their entire curated dataset of 100 prompts.
  • SimpleQA (wei2024simpleQA): A dataset of short, fact-seeking queries across diverse topics. The prompts were created to be challenging for frontier models (e.g. GPT-4o accuracy $40\%$). We use 50 randomly-sampled prompts.
  • WildBench (linwildbench): A subset of the WildChat dataset (zhaowildchat). WildChat is a corpus of 1 million user-ChatGPT conversations. WildBench is a filtered subset of WildChat such that tasks are diverse and challenging for models. We use 50 randomly-sampled prompts from the WildBench-V2.
@@ 6.3 Task Classification Into Our Taxonomy
When calculating functional diversity, we use ground-truth task categories for each prompt based on the source dataset and human-annotation. When generating responses in our task-anchored sampling technique, we use the model’s task categorization of the prompt.
Table 6 shows the number of prompts by ground-truth category and dataset. For the SimpleQA and MATH-500 datasets, we classify prompts as category A (well-specified objective) and category D (problem-solving objective), respectively. For Community Alignment, NoveltyBench, and WildBench, two authors independently classified these prompts into categories. 11 prompts received disagreeing labels, which the annotators resolved after discussion. 6 prompts did not fit into our taxonomy (all from WildBench) due to missing information (e.g. prompts that referenced an unattached document) or language (we only evaluate English prompts).
We determine models’ task categorization of each prompt using the judge prompt below. The agreement rate with ground-truth categories is 82% for GPT-4o, 86% for Claude-4-Sonnet, 84% for Gemini-2.5-Flash, 56% for Llama-3.1-8B-Instruct, and 46% for Mistral-7B-Instruct-v0.3.
Table 6: Number of prompts per dataset and taxonomy category.
Community Math 500 MacGyver Novelty Bench Simple QA Wild Bench Total
Alignment
Well-Specified Objective 2 0 0 1 50 0 53
Underspecified Objective 0 0 0 16 0 0 16
Random Generation 0 0 0 14 0 0 14
Problem-Solving Objective 0 50 0 0 0 5 55
Problem-Solving Subjective 0 0 50 0 0 0 50
Encyclopedia Inquiry 9 0 0 2 0 12 23
Creative Writing 1 0 0 23 0 21 45
Advice or Opinions 38 0 0 44 0 6 88
Total 50 50 50 100 50 44 344
Read the prompt below and decide which task category it belongs to. For prompts that have objective responses, choose from categories A, B, C, or D. For prompts that have subjective responses, choose from categories E, F, G, or H.
@@
E - Problem Solving Subjective: Task to generate an answer with reasoning or explanations for a problem with many verifiably correct answers.
F - Encyclopedia Inquiry Subjective: Task to generate information about real-world societies, traditions, events, or social domains, where there are credible references.
G - Creative Generation Subjective: Task to generate a response that involves creative expression where there are potentially infinite subjective responses.
H - Advice or Opinion Subjective: Task to generate a response that gives advice, opinions, or feedback on specific topics or scenarios.
For the prompt above, only output the assigned task category (A, B, C, D, E, F, G, or H) without any additional text.
(a) GPT-4o
(b) Claude-4-Sonnet
(c) Gemini-2.5-Flash
(d) Llama-3.1-8B-Instruct
(e) Mistral-7B-Instruct-v0.3
Figure 4: Heatmaps showing recall for models’ task classification (proportion of prompts classified by the model into each task category, conditioned on each ground-truth task category).
@@ 6.4 Sampling Strategies
We generate 5 responses per prompt using temperature sampling, in-context regeneration, and system prompt sampling. For both in-context regeneration and system prompt sampling, we always use temperature 1.0 for GPT-4o and Gemini-2.5-Flash, and temperature 0.5 for Claude-4-Sonnet. For in-context regeneration, the first response is generated with the original prompt. Subsequent responses are generated using the regeneration prompts below. For system prompt sampling, multiple responses are extracted from a single generation. The model is instructed to produce multiple responses separated by a delimiter, allowing them to be de-aggregated with regular expressions. For task-anchored approaches, the task category for the prompt is determined by the model’s self-categorization of the task (c.f. Appendix 6.3). For all sampling strategies, we set the nucleus sampling parameter to 0.9 and the maximum number of output tokens to 1024. While we generate responses iteratively, parallel inference techniques could help scale our approach dong2025generalized.
@@
We use the following prompt for the general system prompting approach. This prompt is similar to the one used in zhang2025cultivating, except they use the phrase “Generate {num_responses} that represent diverse values”. We generalize this instruction to apply to all the categories in our taxonomy, since zhang2025cultivating focus on more subjective tasks.
@@
zhang2025noveltybench use the following prompt for in-context regeneration. We call this a “general” prompt because there is no task dependence.
@@
Problem-Solving Subjective The following problem has multiple correct answers, and may be solved using different problem-solving strategies. Generate {num_responses} different solutions, each with a different answer or problem-solving strategy. Can you solve the problem using a different strategy? The problem has multiple correct answers, and may be solved using different problem-solving strategies.
@@ 6.5 Measuring Functional Diversity
We validate the LLM-judges on a stratified random sample of 225 response pairs across models, tasks, and sampling strategies. Two authors independently labeled these prompts for functional diversity. Table 8 reports the inter-rater agreement between annotators and each LLM-judge. When compared to the ensemble of LLM-judges, annotator 1 achieved 77.7% agreement and annotator 2 achieved 79.0% agreement. This is comparable to zhang2025noveltybench’s evaluation of functional diversity in creative tasks, where a fine-tuned judge achieved 79% agreement with human annotators.
@@
To compare with functional diversity, we also measure diversity using three general diversity metrics that do not have task-dependence. For embedding diversity, we generate response embeddings using the gemini-embedding-001 model (with 3072-dimensional embeddings). For compression diversity, we use gZip to compress the concatenated text of response pairs shaib2024standardizing.
Definition 6.1 (Vocabulary Diversity).
Definition 6.2 (Embedding Diversity).
$d_{\mathrm{embed}}(y_{a},y_{b}):=1-\cos(e(y_{a}),e(y_{b})),$
Definition 6.3 (Compression Diversity).
Given two responses $y_{a},y_{b}\in\mathcal{Y}$, the compression diversity between $y_{a}$ and $y_{b}$ is
$d_{\mathrm{compression}}(y_{a},y_{b}):=\frac{\text{compressed size of }y_{a}||y_{b}}{\text{size of }y_{a}||y_{b}},$
where $y_{a}||y_{b}$ is the concatenated text of responses $y_{a}$ and $y_{b}$.
Table 8: Annotator Agreement
@@
Category Task-Anchored Functional Diversity Def. Diversity Judge Options
@@
Problem-Solving Subjective The prompt involves solving a problem with multiple correct answers, and may be solved using different problem-solving strategies. Responses are functionally equivalent if they represent the same answer and problem-solving strategy. Output YES if the responses represent the same answer and problem-solving strategy. Output NO if the responses represent different answers or problem-solving strategies.
@@
Table 10: Examples of Functionally Diverse Responses by Category
Category Prompt Response 1 Response 2
Well-Specified
Objective
What is the name of the project that won the 2014 ACM Software System Award? The project that won the **2014 ACM Software System Award** is **Apache Hadoop**. Apache Hadoop is an open-source framework for distributed storage and processing of large datasets, which has had a significant impact on big data processing and analytics. The project that won the 2014 ACM Software System Award is **The Eclipse Project**.
Underspecified
Objective
Name a Harry Potter Book. Harry Potter and the Sorcerer’s Stone Harry Potter and the Half-Blood Prince
Random
Generation
Pick a number between 1 and 100. Now tell me what you picked. I picked the number 42. I picked the number 87.
Problem-Solving
Objective
If $2^{8}=4^{x}$, what is the value of $x$? We know that $4^{x}=(2^{2})^{x}$. Using the property of exponents $(a^{m})^{n}=a^{m\cdot n}$, this becomes: $4^{x}=2^{2x}.$ The original equation $2^{8}=4^{x}$ now becomes: $2^{8}=2^{2x}.$ Since the bases are the same, we equate the exponents: $8=2x.$ Solve for $x$ by dividing both sides by 2: $x=4.$ We start with the equation $2^{8}=4^{x}$. Take the natural logarithm (ln) of both sides: $\ln(2^{8})=\ln(4^{x}).$ Using the logarithmic rule $\ln(a^{b})=b\cdot\ln(a)$, this becomes: $8\cdot\ln(2)=x\cdot\ln(4).$ Express $\ln(4)$ as $\ln(2^{2})$, which simplifies to $2\cdot\ln(2)$. Substituting this back: $8\cdot\ln(2)=x\cdot(2\cdot\ln(2)).$ Cancel $\ln(2)$ from both sides: $8=2x.$ Solve for $x$: $x=4.$
Problem-Solving
Subjective
You need to tighten a loose door knob but you don’t have a screwdriver. The available items are a round-tipped hairpin, a brittle plastic fork, a flimsy nail file, a pair of rusted pliers, a thin rubber band, and a butter knife. The door knob screws are not the standard slot or Phillips head, but are special star head screws which cannot be tightened using a butter knife or hairpin due to the unusual shape and the risk of stripping the screw head. How can you tighten the door knob? You can use the pair of rusted pliers as a workaround tool. Although they are rusted, they might still have enough grip to grab onto the edges of the star head screw. Position the pliers carefully around the head of the screw and twist it in the tightening direction. Apply steady pressure and try to avoid slipping. Clean the rust off the pliers beforehand, if possible, to improve their grip. Since none of the tools are directly suitable for a star head screw, you can try securing the loose door knob in place temporarily. Use the thin rubber band to wrap tightly around the door knob base, compressing it against the door. This won’t tighten the screws but can stabilize the knob until you acquire the correct star head screwdriver for a permanent fix.
Encyclopedia
Inquiry
Did any major countries reinstate the death penalty after prohibiting it? While it is not common, some nations have reinstated the death penalty after abolishing it. For example, Pakistan imposed a moratorium on executions in 2008, effectively halting the use of the death penalty. However, in 2014, the country lifted the moratorium following a devastating terrorist attack, and executions resumed. This demonstrates how security concerns can influence such decisions. Historically, a few countries have reinstated the death penalty after initially banning or suspending it. The United States offers an example at the state level, where capital punishment was temporarily halted in the 1970s after the Supreme Court’s Furman v. Georgia decision but reinstated later following the Gregg v. Georgia ruling in 1976. Nationally, reinstatements are usually tied to shifting societal attitudes or judicial rulings.
Creative
Writing
Write a short love poem with 4 lines. Your laughter blooms like spring’s embrace, A gentle warmth I cannot replace. In your eyes, a world so true, Forever, my heart beats for you. The ocean hums your name to me, Its waves reflect eternity. My soul is bound, your love’s my art, Forever etched within my heart.
Advice or
Opinions
I am naming my Indian son what one name should I name him? I get anxiety when given multiple options so please only give me one name. Consider naming your son “Kabir.” This name is associated with wisdom and spirituality, inspired by the renowned Indian poet and saint. You might like the name “Rohan.” It is simple, easy to pronounce globally, and represents growth and ascent.
Table 11: Examples of Homogeneous Responses by Category
Category Prompt Response 1 Response 2
Well-Specified
Objective
What is the name of the project that won the 2014 ACM Software System Award? The project that won the **2014 ACM Software System Award** is **The Coq Proof Assistant**. Coq is a formal proof management system, providing a formal language to write mathematical definitions, executable algorithms, and theorems, together with tools to verify the correctness of these proofs. The project that won the 2014 ACM Software System Award is **Coq**, a formal proof management system. Coq is widely used in the field of computer science for formalizing mathematical proofs and for developing software with guarantees of correctness.
Underspecified
Objective
Name a Harry Potter Book. Harry Potter and the Sorcerer’s Stone Harry Potter and the Sorcerer’s Stone (or Harry Potter and the Philosopher’s Stone in the UK).
Random
Generation
Pick a number between 1 and 100. Now tell me what you picked. Sure! I’ll pick the number **42**. Sure! I pick 42.
Problem-Solving
Objective
If $2^{8}=4^{x}$, what is the value of $x$? We know that $4^{x}=(2^{2})^{x}$. Using the property of exponents $(a^{m})^{n}=a^{m\cdot n}$, this becomes: $4^{x}=2^{2x}.$ The original equation $2^{8}=4^{x}$ now becomes: $2^{8}=2^{2x}.$ Since the bases are the same, we equate the exponents: $8=2x.$ Solve for $x$ by dividing both sides by 2: $x=4.$ Step 1: Express $4^{x}$ in terms of base 2 Since $4=2^{2}$, we can rewrite $4^{x}$ as: $4^{x}=(2^{2})^{x}$ Using the property of exponents $(a^{m})^{n}=a^{m\cdot n}$, this becomes: $4^{x}=2^{2x}$ Step 2: Substitute into the equation The equation now becomes: $2^{8}=2^{2x}$ Step 3: Equating exponents Since the bases are the same ($2$), we can equate the exponents: $8=2x$ Step 4: Solve for $x$ Divide both sides of the equation by 2: $x=\frac{8}{2}=4$ Final Answer: $\boxed{4}$
Problem-Solving
Subjective
You need to tighten a loose door knob but you don’t have a screwdriver. The available items are a round-tipped hairpin, a brittle plastic fork, a flimsy nail file, a pair of rusted pliers, a thin rubber band, and a butter knife. The door knob screws are not the standard slot or Phillips head, but are special star head screws which cannot be tightened using a butter knife or hairpin due to the unusual shape and the risk of stripping the screw head. How can you tighten the door knob? Wrap the thin rubber band around the head of the screw to improve grip. Use the tips of the rusted pliers to carefully turn the screw, pressing firmly to ensure the pliers engage with the rubber band and the screw head without slipping. Tighten the screw gradually while maintaining steady pressure. Use the rusted pliers to grip the edges of the star head screw. Apply steady pressure and carefully turn the pliers to tighten the screw. If additional grip is needed, place the thin rubber band between the screw head and the pliers to increase friction and prevent slipping.
Encyclopedia
Inquiry
Did any major countries reinstate the death penalty after prohibiting it? The Philippines reinstated the death penalty in 1993 after having abolished it in 1987, before abolishing it again in 2006. Pakistan reinstated executions in 2014 after a six-year moratorium, primarily in response to terrorism. Sri Lanka has made moves toward resuming executions for drug-related offenses despite not carrying any out since 1976. In the United States, individual states like New York have reinstated the death penalty after previously prohibiting it, though some of these efforts were later reversed or deemed unconstitutional. The Philippines reinstated the death penalty in 1993 after abolishing it in 1987, but it was abolished again in 2006. Pakistan lifted a six-year moratorium on the death penalty in 2014 and resumed executions, citing security concerns. In the United States, some states reinstated the death penalty after earlier bans, such as New York in the 1990s, though in some cases, these reinstatements were later overturned. Sri Lanka has expressed intentions to resume executions despite not carrying out any since 1976.
Creative
Writing
Write a short love poem with 4 lines. Beneath the stars, your gaze feels near, A whispered promise, soft and clear. In your embrace, my heart takes flight, You are my day, my moon, my night. Beneath the stars, your eyes take flight, A gentle glow, my endless light. With every breath, my heart you keep, A love eternal, vast and deep.
Advice or
Opinions
I am naming my Indian son what one name should I name him? I get anxiety when given multiple options so please only give me one name. Sure! A beautiful and meaningful Indian name for your son could be **Arjun**. It is rooted in Indian heritage, symbolizes bravery and focus, and has a timeless charm. How about the name **Arjun**? It is a strong, classic Indian name with roots in the Mahabharata, symbolizing bravery, focus, and virtue.